OpenStudy (anonymous):

Simplify PLEASE? 56 x 4

OpenStudy (anonymous):

http://symbolab.com/solver/inequalities-calculator This may help. ;) (It isn't only for inequalities)

OpenStudy (anonymous):

\[\frac{ 7xy }{ x }+\frac{ 10y^2 }{ 5y } = 7y + 2y = 9\] That is a very simple function. You can solve it by yoursef. Just try a little bit. It is solved like this: x's in the first term cancels each other(7y), y's in the second term cancels each other(10y/5). 10y/5 = 2y 7y+2y = 9y

OpenStudy (anonymous):

can give you one more?